how to solve this equation and check the solution
-2x + 5 = -79
-3x +19 = -8
Please go step by step explaining how you got your result.

Answer by Alan3354(69443)

how to solve this equation and check the solution
-2x + 5 = -79
-3x +19 = -8
---------
There are 2 equations.
-2x + 5 = -79
Subtract 5
-2x = -84
Divide by 2
-x = -42
Multiply by -1
x = 42
--------------
Do the other one the same way.